Hi,i'm using factoring to solve problems,and I can't figure this out. Here's the question: A rectangular garden 30 meters by 40 meters has two paths of equal width crossing through it.Find the width of each path if the total area covered by the paths is 325m^2.Thanks a lot!

Answer by Nate(3500)

Hello,
Your Question:
Hi,i'm using factoring to solve problems,and I can't figure this out. Here's the question: A rectangular garden 30 meters by 40 meters has two paths of equal width crossing through it.Find the width of each path if the total area covered by the paths is 325m^2.Thanks a lot!
-------------------------------------------------------------------------------
My Answer: (Imagine this as the garden's shape: [] is the boundary of the garden; (I) being the path; --- being whatever is in the garden)
[-I-I-]30
40
We know that it is 30m by 40m. If the two paths are straight, then we know that the length of the path going up and down is 30m. So you could use the formula for a rectangle to solve this, and since it is two rectangles, you would multiply by two (2LW)!
2LW=325 (divide by two)
LW=325/2
(we know that the length of the path going along the width is 30m so, we plug in 30m for width)
L(30)=325/2 (divide by 30)
L=325/60
So the width of the path going from left to right would be 325/60 or 65/12 (simplified)!